Find the minimum sample size n needed to estimate u for the given values of c, o, and E.
c=0.90
o=9.3
E=2
prelimary sample of 30 members--> n=___

Answer by Boreal(15235)

error=t(0.95,df=?)*s/sqrt(n); I will use s here.
2=1.699*9.3/sqrt(n)
square both sides, multiply by n and divide by 4, the 2^2
n=15.80^2/4
=62.41, so use 63
t(0.95, df=62) is 1.670
2=1.67*9.3/sqrt(62)
the right side=1.956
try sample size 61 to see if that works. it does, with the quotient < 2.
n=60 gives error of 2.005, sho the minimum sample size is 61.